Big Rig Delivery Service has the following plant assets: Communications Equipment: Cost, $3,840 with useful life of 8 years; Delivery Equipment: Cost, $17,568 with useful life of 12 years; and Computer: Cost, $13,440 with useful life of 4 years. Assume the residual value of all the assets is zero and the straight-line method is used.

Big Rig's monthly depreciation journal entry will include a ________.

A. credit to Depreciation Expense - Equipment of $5,304

B. debit to Depreciation Expense - Equipment of $5,304

C. debit to Accumulated Depreciation - Equipment of $442

D. credit to Accumulated Depreciation - Equipment of $442
Answer #1

Total annual depreciation

= (3840/8)+(17568/12)+(13440/4)

= 480+1464+3360

= 5304

Depreciation expense per month = 5304/12 months

= 442

Option D is the answer